BexleyBrooks:

I need some help with these math questions lol 1. x+3+7=2(x)-10+x What is X??? 2. 3(5-3(x)+7(x)=9 What is X??

D14610:

For the first equation, let's simplify and solve for : x+3+7= 2x - 10 + x This simplifies to: x + 10 = 3x - 10 Rearranging terms gives: 20 = 2x So, a = 10 for the first equation. For the second equation, let's simplify and solve for : 3(5-3x+7x)=9 This simplifies to: 3(5+4x) = 9 15 + 12x = 9 12x = -6 So, x=-0.5 for the second equation.

Phantomdex:

For 1: x+10=3x−10 10=3x−x−10 10=2x−10 10+10=2x−10+10 20=2x 20/2 = 2x/2 10=x So, x=10 For 2: 3(5−3x+7x)=9 3(5+4x)=9 15+12x=9 15+12x−15=9−15 12x=−6 12x/12 = -6/12 x = -1/2